Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 583
Entry time: 29 Jan 2025, 09:50
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 29.01.2025 |
|
|
General Announcement:
- Steffen Hauf comments that DOC is active from today and also requests SXP and SCS to get in touch with DRC on needed nodes.
- UOM will restart next Friday at 9AM.
CTRL:
- Successful bring-up days were concluded.
DA:
- One last bring-up day to be done today.
EEE:
- Successful bring-up days.
VAC:
- All fine with VAC.
XPD:
- Concluded last action of XPD preparations. Small issues were addressed.
- ZZ in XTD6 yesterday: Shimadzu camera ZZ everything is fine by restarting PC.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- ZZ in XTD10: stuck pre-mirror in soft-mono due to Elmo controller; physical activity brought back.
- 1mm of limit switch of M5 SA3 is also corrected.
- ZZ in SA2: piezo cables of M1 was swapped and working back.
- Bring up days went fine yesterday.
LAS:
- Nothing to report.
UNSYS:
- Nothing to report.
SPB/SFX:
- PBLM was found in front of the beam; took a time to realize that, fixed by now.
FXE:
- Nothing to report.
MID:
- Issues with power slits at tunnel; but fix ongoing; not a show stopper.
SCS:
- Nothing to report.
SQS:
- Not present.
SXP:
- Getting beam today. |